the probability that the operations being successful is 3/4 . what is the probability of winning 3 consecutive operations?
Answer by Theo(13342)

p(success) = 3/4

p(success three times in a row) = 3/4 * 3/4 * 3/4 = 27/64

the probability of 3 successes 1 time in a row is 3/4.
the probability of 3 successes 2 times in a row is 3/4 * 3/4.
the probability of 3 successes 3 times in a row is 3/4 * 3/4 * 3/4.